16. An aqueous solution has a normal boiling point of 103.5°C. What is the freezing point of this solution? For water and
An aqueous solution has a normal boiling point of 103.5°C. What is the freezing point of this solution? For water and
-3.6°C
-3.5°C
-13 °C
-0.96°C
SubmitMy An
1
Expert's answer
2017-09-08T07:06:15-0400
Solution: Find the molality of the solution: ΔT = imKb ΔT = 103.5 - 100 = 3.5 oC i = 1 m = molality Kb = boiling point constant = 0.51 m = (3.5)/(1)(0.51) = 6.86 m Now, we find the freezing point of this solution ΔT = - imKf ΔT = - (1)(6.86)(1.86) ΔT = - 12.8 oC ≈ - 13 oC
Answer: - 13 oC